# Shuai Wang, Xin Zhong

### v1

@@ 1.2 Main result
$\frac{3\mu}{3\mu+\lambda}\leq\min\bigg\{\frac{1}{2\Lambda(4)},a^{-1}\Lambda^{-1}\bigg(\frac{25\alpha+11}{\alpha-1}\bigg)\bigg\},$ (1.12)
@@
$\begin{cases}0\leq\rho\in L^{\infty}(0,T;L^{\theta}(\mathbb{R}^{2}_{+}))\cap C([0,T];L^{q}(\mathbb{R}^{2}_{+})),\ \ \text{for any}\ q\in[1,\theta),\\ \bar{x}^{\alpha}\rho\in L^{\infty}(0,T;L^{1}(\mathbb{R}^{2}_{+})),\ \ \rho\mathbf{u}\in C([0,T];H^{-1}(\mathbb{R}^{2}_{+})),\\ \nabla\mathbf{u}\in L^{2}(\mathbb{R}^{2}_{+}\times(0,T)),\ \ \big(\sigma^{\frac{1}{2}}\sqrt{\rho}\dot{\mathbf{u}},\ \sigma^{\frac{3}{2}}\nabla\dot{\mathbf{u}}\big)\in L^{2}(\mathbb{R}^{2}_{+}\times(0,T)),\end{cases}$ (1.14)
@@
The present work improves the result of Perepelitsa [31]. On the one hand, the presence of vacuum is allowed both at interior and infinity. On the other hand, in contrast to the $L^{\infty}$ density assumption in [31], our solutions are constructed within an $L^{\theta}$ density framework, which permits unbounded densities and enables a reduction of the time growth in the far-field analysis through the relaxed spatial integrability of the pressure in the sense that $P(\rho)\in L^{q}$ for $1<q=\theta/\gamma<\infty$.
@@
The index $\theta$ in (1.8) is a technical parameter determined by weighted inequalities in Lemma 2.4 and large-time analysis in (3), and may be relaxed by delicate analysis; see Subsection 1.3 for details. In addition, it should be mentioned that it is not clear whether the regularity assumption on the initial velocity in Theorem 1.1 can be generalized to $H^{s}$ for some $0<s<1$, which has been achieved in the whole plane case [14] as the initial density is bounded from above and below (among other conditions).
@@ 1.3 Strategy of the proof
We now outline the main ideas and difficulties in the proof. The argument relies on the construction of global smooth approximate solutions via a delicate limiting procedure. This approach consists of two fundamental steps. First, we apply the local existence theory for initial densities that are strictly positive (Lemma 2.1). Second, we pass to the limit as the lower bound of the initial density tends to zero (see Section 4 for details). Bridging these two steps crucially depends on deriving suitable a priori estimates that are uniform with respect to the lower bound of the initial density and ensure the desired regularity.
@@
Moreover, obtaining any decay control at infinity within the weak solution framework is particularly delicate and challenging, especially in the presence of Dirichlet boundary conditions. This is due to the weakness of the available a priori bounds, such as the crude control of $\|\nabla\mathbf{u}\|_{L^{q}}$, and to the lack of higher-order estimates, which leaves no mechanism to derive decay in this setting. In view of the tools at hand, the only information we can employ is the $L^{\theta}$ bound concerning the density. As an attempt, we exploit the material derivative weighted by the density (see (3.56) and (3.59)), and obtain only a very poor growth bound:
@@
Unlike in the strong-solution setting, the route based on the decay-control for higher derivatives of the velocity field is no longer available, so the decay mechanism in our setting leads us to re-examine the density and the pressure $P(\rho)$. By exploiting the higher $L^{60}$ integrability of the pressure together with the spatial embedding in two-dimensional setting, we sharpen the above estimate and reduce the time growth (see (3)):
@@
which can now be enclosed. Furthermore, this integrability requirement together with Lemma 2.4 dictates our choice of the index
$\theta=\max\big\{4(2\alpha+1)/(\alpha-1),60\gamma\big\}.$
As a heuristic indication for a possible refinement of this index, we reconsider the growth bounds (2.1) and (3.56) with $\|\rho\dot{\mathbf{u}}\|_{L^{r}}\lesssim(1+t)^{4}$ for $2<r<\infty$. To accommodate the far-field behavior, the choice of the index is guided by the Gagliardo–Nirenberg inequality (see Lemma 2.2) with
@@
$\tilde{\theta}=\max\big\{\tilde{\zeta}(\alpha,r),q\gamma=30\gamma r/(r-2)\big\}>30\gamma,$
where we take $r=4$ in the present work.
@@
becomes difficult to control analytically for $\sigma=\sigma(t)\triangleq\min\{1,t\}$, particularly in view of the lack of a satisfactory bound for $\|\nabla\mathbf{u}\|_{L^{4}}$. To overcome this obstacle, we split the time interval into two regimes and carefully isolate the singular behavior of the pressure term near the initial time (see (3) and (3.20)). This procedure allows us to focus on
@@
which can be enclosed by the estimates derived in (3). In a certain sense, the interplay between reduced density regularity and the Dirichlet boundary condition manifests itself in stronger temporal singularities near the initial time.
Having resolved the above difficulties, we further apply Zlotnik’s inequality (see Lemma 2.5) to obtain a time-uniform upper bound for $\|\rho\|_{L^{\theta}}$. A key step in this procedure is the control of $\|\widetilde{F}\|_{L^{\infty}}$, as shown in (3) and (3), where the density-weighted inequality established in Lemma 2.4 and the spatial weighted estimate in Lemma 3.5 play central roles. With these a priori estimates in hand, we obtain the desired regularity for the density, which is higher than that of Lions–Feireisl’s class. At last, by a standard compactness argument combined with the continuity method, we construct a global weak solution in the sense of Definition 1.1 that satisfies (1.14).
@@ 2.2 An exposition on the boundary
Let $(\rho,\mathbf{u})$ be a smooth solution to the problem (1.1)–(1.4). For any $1<q<\infty$, there exists a constant $C>0$ depending only on $\mu$, $\lambda$, and $q$ such that
@@ 3 A priori estimates
First of all, if we multiply (1.1)1 by a cut-off function and use a standard limit procedure, one can obtain that
@@
$\leq CC_{0}+C\int_{0}^{T}\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma^{\prime}P|\nabla\mathbf{u}|\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$
@@
$\leq\int_{\sigma(T)}^{T}\|P\|_{L^{4}}\|\nabla\mathbf{u}\|_{L^{2}}\big(\|\sqrt{\rho}\|_{L^{4}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{4}}\big)\mathrm{d}t$
@@
$A_{1}(T)+A_{2}(T)\leq CC_{0}^{\frac{5\alpha+4}{23\alpha+13}}+C\int_{0}^{T}\sigma\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$ (3.30)
@@
After multiplying it by $3\sigma^{3}P^{2}$ and integration by parts, we obtain from Young’s inequality that
@@
$\leq CK+C\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t+C\int_{0}^{\sigma(T)}\sigma\|P\|_{L^{4}}^{4}\mathrm{d}t.$ (3.44)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}$ $\leq\frac{\theta-1}{2\mu+\lambda}\int\rho^{\theta}|\widetilde{F}|\mathrm{d}\mathbf{x}+\frac{2\mu(\theta-1)}{(3\mu+\lambda)(2\mu+\lambda)}\int\rho^{\theta}\big|\mathcal{Q}_{3}(P)\big|\mathrm{d}\mathbf{x}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}\leq-\frac{\theta-1}{3(2\mu+\lambda)}\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}}.$ (3.53)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}\leq-\frac{\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\big(\|\rho\|_{L^{\theta}}^{\theta}\big)^{\frac{\theta}{\theta-\gamma}}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}},$ (3.54)
@@
$y(t)=\|\rho\|_{L^{\theta}}^{\theta},\ \ f(y)=-\frac{\widetilde{c}(\theta-1)}{3(2\mu+\lambda)}y^{\frac{\theta}{\theta-\gamma}}(t),\ \ b(t)=\int_{0}^{t}\frac{\theta-1}{2\mu+\lambda}\|\widetilde{F}\|_{L^{\infty}}y(\tau)\mathrm{d}\tau,$
one deduces from (3.54) that
@@
Applying Lemma 2.5 with
@@
then it holds that
$f(\xi)=-\frac{\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\xi^{\frac{\theta}{\theta-\gamma}}\leq-N_{1}=0,\ \ \ \text{for all}\ \xi\geq\xi^{*}=1.$
Moreover, one has that
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int t\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+\mu t\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\leq C\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+Ct\|P\|_{L^{4}}^{4}+Ct\|\nabla\mathbf{u}\|_{L^{4}}^{4}.$
@@
which together with (3.56), Lemma 2.8, Young’s inequality, and (3.28) that
@@
$N_{0}=C_{5}(\hat{\rho})C_{0}^{\frac{5\alpha+4}{23\alpha+13}},\ \ N_{1}=\widetilde{C},\ \ \xi^{*}=\frac{3\widetilde{C}(2\mu+\lambda)}{\tilde{c}(\theta-1)}+1,$
@@
$f(\xi)=-\frac{\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\xi^{\frac{\theta}{\theta-\gamma}}\leq-N_{1},\ \ \ \text{for all}\ \xi\geq\xi^{*}.$
Furthermore, it holds that
@@
$\bar{x}^{\alpha}\rho_{0}^{\epsilon}\in L^{1},\ \ (\rho_{0}^{\epsilon}-\phi\epsilon)\in H^{2},\ \ \inf_{\mathbf{x}\in\widetilde{B}_{1/\epsilon}}\{\rho_{0}^{\epsilon}(\mathbf{x})\}\geq\epsilon,\ \ \mathbf{u}_{0}^{\epsilon}\in D^{2}\cap D_{0}^{1},\ \sqrt{\rho_{0}^{\epsilon}}\mathbf{u}_{0}^{\epsilon}\in L^{2},$
@@
$\sup_{t\in[0,T]}\|\bar{x}^{\alpha}\rho^{\epsilon}\|_{L^{1}}\leq C(T).$
@@
Hence, by the Aubin–Lions lemma, up to the extraction of a subsequence (still labeled by $\epsilon$), we have

### v2

@@ 1.2 Main result
$\frac{3\mu}{3\mu+\lambda}\leq\min\bigg\{\frac{1}{2\Lambda(4)},\Lambda^{-1}\bigg(\frac{25\alpha+11}{\alpha-1}\bigg)\bigg\},$ (1.12)
@@
$\begin{cases}0\leq\rho\in L^{\infty}(0,T;L^{\theta}(\mathbb{R}^{2}_{+}))\cap C([0,T];L^{q}(\mathbb{R}^{2}_{+})),\ \ \text{for any}\ q\in[1,\theta),\\ \bar{x}^{\alpha}\rho\in L^{\infty}(0,T;L^{1}(\mathbb{R}^{2}_{+})),\ \ \sqrt{\rho}\mathbf{u}\in C([0,T];L^{2}(\mathbb{R}^{2}_{+})),\ \ \nabla\mathbf{u}\in L^{2}(\mathbb{R}^{2}_{+}\times(0,T)),\\ \big(\sigma^{\frac{1}{2}}\nabla\mathbf{u},\ \sigma^{\frac{3}{2}}\sqrt{\rho}\dot{\mathbf{u}}\big)\in L^{\infty}(0,T;L^{2}(\mathbb{R}^{2}_{+})),\ \ \big(\sigma^{\frac{1}{2}}\sqrt{\rho}\dot{\mathbf{u}},\ \sigma^{\frac{3}{2}}\nabla\dot{\mathbf{u}}\big)\in L^{2}(\mathbb{R}^{2}_{+}\times(0,T)),\end{cases}$ (1.14)
@@
The present work improves the result of Perepelitsa [31]. On the one hand, the presence of vacuum is allowed both at interior and infinity. On the other hand, in contrast to the $L^{\infty}$ density assumption in [31], our solutions are constructed within an $L^{\theta}$ density framework, which permits unbounded densities and captures the far-field behavior while mitigating the time growth in the far-field analysis, owing to the relaxed spatial integrability of the pressure, in the sense that $P(\rho)\in L^{q}$ for $1<q=\theta/\gamma<\infty$.
@@
The index $\theta$ in (1.8) is a technical parameter determined by weighted inequalities in Lemma 2.4 and large-time analysis in (3), and may be relaxed by delicate analysis; see Subsection 1.3 for details. The assumption (1.12) on viscosity is introduced solely to handle the difficulties arising from the no-slip boundary. In addition, it should be mentioned that it is not clear whether the regularity assumption on the initial velocity in Theorem 1.1 can be generalized to $H^{s}$ for some $0<s<1$, which has been achieved in the whole plane case [14] as the initial density is bounded from above and below (among other conditions).
@@ 1.3 Strategy of the proof
We now outline the main ideas and difficulties in the proof. The argument relies on the construction of global smooth approximate solutions via a delicate limiting procedure. This approach consists of two fundamental steps. First, we apply the local existence theory for initial densities that are strictly positive (Lemma 2.1). Second, we pass to the limit as the lower bound of the initial density tends to zero (see Section 4 for details). Bridging these two steps crucially depends on deriving suitable a priori estimates that are uniform with respect to this lower bound and ensure the desired regularity.
@@
Although these less tractable estimates do not close on their own, they allow the argument to proceed and ultimately close once the small-energy assumption is fully exploited.
Moreover, obtaining any decay control at infinity within the weak solution framework is particularly delicate and challenging, especially in the presence of Dirichlet boundary conditions. This is due to the limited usefulness of the available a priori bounds, such as the crude control of $\|\nabla\mathbf{u}\|_{L^{q}}$, and to the lack of higher-order estimates, which leaves no mechanism to derive decay in this setting. In view of the tools at hand, the only information we can employ is the $L^{\theta}$ bound concerning the density. As an attempt, we exploit the material derivative weighted by the density (see (3.56) and (3.59)), and obtain a polynomial growth bound:
@@
Unlike in the strong-solution setting, the route based on the decay-control for higher derivatives of the velocity field is no longer available, so we abandon the decay-based approach and re-examine the density and the pressure $P(\rho)$. By exploiting the integrability of the pressure together with the spatial embedding in two-dimensional setting, we sharpen the above estimate and reduce the time growth (see (3)):
@@
which can now be enclosed under (1.13). Furthermore, this integrability requirement together with Lemma 2.4 dictates our choice of the index
$\theta\geq\max\big\{4(2\alpha+1)/(\alpha-1),60\gamma\big\}.$
As a heuristic indication for a possible refinement of this index, we revisit the growth bounds in (3.56) and (2.1) with $\varrho\in L^{1}\cap L^{\tilde{\zeta}(\alpha,r)}$, which yield $\|\rho\dot{\mathbf{u}}\|_{L^{r}}\lesssim(1+t)^{4}$ for $2<r<\infty$. To capture the far-field behavior, the choice of the index is guided by the Gagliardo–Nirenberg inequality (see Lemma 2.2) with
@@
$\tilde{\theta}\geq\max\big\{\tilde{\zeta}(\alpha,r),q\gamma=30\gamma r/(r-2)\big\}>30\gamma,$
where we take $r=4$ and $\tilde{\zeta}(\alpha,r)=4(2\alpha+1)/(\alpha-1)$ in the present work.
@@
becomes difficult to control analytically for $\sigma=\sigma(t)\triangleq\min\{1,t\}$, particularly in view of the lack of a satisfactory bound for $\|\nabla\mathbf{u}\|_{L^{4}}$. To overcome this obstacle, we split the time interval into two regimes and carefully account for the initial-time singularity associated with the velocity gradient by means of an $L^{6}$ estimate (see (3) and (3.20)). This procedure allows us to focus on
@@
which can be enclosed by the estimates derived in (3) once the effective viscous flux has been controlled. In a certain sense, the reduced density regularity manifests itself in stronger temporal singularities near the initial time.
Having resolved the above difficulties, we further invoke Zlotnik’s inequality (see Lemma 2.5) to obtain a time-uniform upper bound for $\|\rho\|_{L^{\theta}}$. A key step in this procedure is the global control of $\|\widetilde{F}\|_{L^{\infty}}$, as shown in (3) and (3), where the density-weighted inequality established in Lemma 2.4 and the spatial weighted estimate in Lemma 3.5 play central roles. With these a priori estimates in hand, we obtain the desired regularity for the density, in particular a level of integrability stronger than that typically available in the Lions–Feireisl’s theory. At last, by a standard compactness argument combined with the continuity method, we construct a global weak solution in the sense of Definition 1.1 that satisfies (1.14).
@@ 2.2 An exposition on the boundary
Let $(\rho,\mathbf{u})$ be a smooth solution to the problem (1.1)–(1.4). For any $2<q<\infty$, there exists a constant $C>0$ depending only on $\mu$, $\lambda$, and $q$ such that
@@ 3 A priori estimates
First of all, if we multiply (1.1)1 by a cut-off function and use a standard limit procedure, then one can obtain that
@@
$\leq C(\hat{\rho})C_{0}^{\frac{22\alpha+14}{23\alpha+13}}+C\int_{0}^{T}\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma^{\prime}P|\nabla\mathbf{u}|\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$
@@
$\leq C\int_{\sigma(T)}^{T}\|P\|_{L^{4}}\|\nabla\mathbf{u}\|_{L^{2}}\big(\|\sqrt{\rho}\|_{L^{4}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{4}}\big)\mathrm{d}t$
@@
$A_{1}(T)+A_{2}(T)\leq CC_{0}^{\frac{2(5\alpha+4)}{23\alpha+13}}+C\int_{0}^{T}\sigma\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$ (3.30)
@@
After multiplying it by $3\sigma^{3}P^{2}$ and integrating by parts, we obtain from Young’s inequality that
@@
$\leq C(K)+C\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t+C\int_{0}^{\sigma(T)}\sigma\|P\|_{L^{4}}^{4}\mathrm{d}t.$ (3.44)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}+\frac{a(\theta-1)}{2\mu+\lambda}\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}$ $\leq\frac{\theta-1}{2\mu+\lambda}\int\rho^{\theta}|\widetilde{F}|\mathrm{d}\mathbf{x}+\frac{2\mu(\theta-1)}{(3\mu+\lambda)(2\mu+\lambda)}\int\rho^{\theta}\big|\mathcal{Q}_{3}(P)\big|\mathrm{d}\mathbf{x}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}\leq-\frac{a(\theta-1)}{3(2\mu+\lambda)}\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}}.$ (3.53)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}\leq-\frac{a\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\big(\|\rho\|_{L^{\theta}}^{\theta}\big)^{\frac{\theta}{\theta-\gamma}}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}},$ (3.54)
@@
$y(t)=\|\rho\|_{L^{\theta}}^{\theta},\ \ f(y)=-\frac{a\widetilde{c}(\theta-1)}{3(2\mu+\lambda)}y^{\frac{\theta}{\theta-\gamma}}(t),\ \ b(t)=\int_{0}^{t}\frac{\theta-1}{2\mu+\lambda}\|\widetilde{F}\|_{L^{\infty}}y(\tau)\mathrm{d}\tau.$
One then deduces from (3.54) that
@@
We now apply Lemma 2.5 with
@@
where one sees that
$f(\xi)=-\frac{a\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\xi^{\frac{\theta}{\theta-\gamma}}\leq-N_{1}=0,\ \ \ \text{for all}\ \xi\geq\xi^{*}=1.$
Therefore, we have that
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int t\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+t\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\leq C\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+Ct\|P\|_{L^{4}}^{4}+Ct\|\nabla\mathbf{u}\|_{L^{4}}^{4}.$
@@
which together with (3.56), Lemma 2.8, Young’s inequality, and (3.28) shows that
@@
$N_{0}=C_{5}(\hat{\rho})C_{0}^{\frac{5\alpha+4}{23\alpha+13}},\ \ N_{1}=\widetilde{C},\ \ \xi^{*}=\frac{3\widetilde{C}(2\mu+\lambda)}{a\tilde{c}(\theta-1)}+1,$
we then deduce that
$f(\xi)=-\frac{a\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\xi^{\frac{\theta}{\theta-\gamma}}\leq-N_{1},\ \ \ \text{for all}\ \xi\geq\xi^{*}.$
@@
$\bar{x}^{\alpha}\rho_{0}^{\epsilon}\in L^{1},\ \ (\rho_{0}^{\epsilon}-\phi\epsilon)\in H^{2},\ \ \inf_{\mathbf{x}\in\widetilde{B}_{1/\epsilon}}\{\rho_{0}^{\epsilon}(\mathbf{x})\}\geq\epsilon,\ \ \mathbf{u}_{0}^{\epsilon}\in D^{2}\cap D_{0}^{1},\ \ \sqrt{\rho_{0}^{\epsilon}}\mathbf{u}_{0}^{\epsilon}\in L^{2}.$
@@
$\sup_{t\in[0,T]}\|\bar{x}^{\alpha}\rho^{\epsilon}(t)\|_{L^{1}}\leq C(T),$
which implies the tightness of $\{\rho^{\epsilon}\}$ at infinity: for any $\delta>0$ there exists $R>0$ such that
[…]
@@
Hence, using (4.3) and the local Aubin–Lions compactness, up to the extraction of a subsequence (still labeled by $\epsilon$), we have